OpenStudy (anonymous):

Which types of triangles can be formed by taking the cross-section of a cube? A. equilateral and isosceles only B. isosceles and scalene only C. equilateral, isosceles, and scalene D. equilateral only

OpenStudy (anonymous):

someone please help !!

OpenStudy (anonymous):

Any clue

OpenStudy (anonymous):

no, i need help.

OpenStudy (anonymous):

So on google look up cross sections of a cube

OpenStudy (anonymous):

Now then see what if one of the answers correlate to a type of triangle listed

OpenStudy (anonymous):

Students will find a way to cut their cubes to form a cross section that is: an equilateral triangle. a triangle that is not equilateral. a parallelogram that is not a rectangle.

OpenStudy (anonymous):

it said that ^

OpenStudy (anonymous):

So it would be d because only equilateral triangle is listed and a cube as equal amount of sides

OpenStudy (anonymous):

thank you so much!!

OpenStudy (anonymous):

Np medal?

OpenStudy (anonymous):

Thx

OpenStudy (anonymous):

np

OpenStudy (anonymous):

Any other question

OpenStudy (anonymous):

If the cube has a side length of 15 inches, what is the surface area of the cube? A. 900 in2 B. 1,350 in2 C. 3,375 in2 D. 90 in2

OpenStudy (anonymous):

So what is the area of a cube look up the formula

OpenStudy (anonymous):

so i go on google and look up what is the area of the cube ? or what is the formula of the cube ?

OpenStudy (anonymous):

Formula

OpenStudy (anonymous):

Volume of a cube = side times side times side. Since each side of a square is the same, it can simply be the length of one side cubed. If a square has one side of 4 inches, the volume would be 4 inches times 4 inches times 4 inches, or 64 cubic inches.

OpenStudy (anonymous):

So plug. In the 15 for a and this should be ur solution A=6a^2=6·15^2=1350

OpenStudy (anonymous):

thanks so much.

OpenStudy (anonymous):

Np